Consider equally risky, all-equity financed firms G and D. Both firms are currently trading at $50 per share. Firm G pays no dividends, but firm D pays all its earnings as dividends. In a year from today, the stock of firm G is expected to be at $65 per share. Firm D is expected to pay a $15 dividend per share at the year-end and its stock price at the year-end is expected to be $50 per share. Capital gains are taxed at half the 40% regular income tax rate, but the dividend income is taxed at 30%.

At what share price should the stock of firm D be trading to-day in a competitive stock market? Show calculations to support your answer.
Answer #1

capital gain from G= (65-50)/50 =.3

after tax gain =.3*(1-.2) =.24 =24%

This should be the return from D from company D as well

after tax return of company D= .24

pre tax return of company D= .24/(1-.3) =.3428

pretax return =dividend/share price

.3428=15/share price

share price of D= 43.76 $